Theory and Applications of Categories, Vol. 24, No. 1, 2010, pp. 1–24.

BICATEGORIES OF SPANS AS CARTESIAN BICATEGORIES

STEPHEN LACK, R.F.C WALTERS, AND R.J. WOOD

Abstract. Bicategories of spans are characterized as cartesian bicategories in which every comonad has an Eilenberg-Moore object and every left adjoint arrow is comonadic.

1. Introduction

Let E be a category with finite limits. For the bicategory Span E, the locally full subbicategory MapSpan E determined by the left adjoint arrows is essentially locally discrete, meaning that each hom category MapSpan E(X, A) is an equivalence relation, and so is equivalent to a discrete category. Indeed, a span x:X [figure] S [figure] A:a has a right adjoint if and only if x:S [figure] X is invertible. The functors

MapSpan E(X, A) [figure] E(X, A) given by (x, a) ↦ ax^{−1}

provide equivalences of categories which are the effects on homs for a biequivalence

MapSpan E [figure] E .

Since E has finite products, MapSpan E has finite products as a bicategory. We refer the reader to [CKWW] for a thorough treatment of bicategories with finite products. Each hom category Span E(X, A) is isomorphic to the slice category E/(X × A) which has binary products given by pullback in E and terminal object 1:X × A [figure] X × A. Thus Span E is a precartesian bicategory in the sense of [CKWW]. The canonical lax monoidal structure

Span E × Span E [figure] Span E [figure] 1

for this precartesian bicategory is seen to have its binary aspect given on arrows by

(X ⟵^{x} S ⟶^{y} A , Y ⟵^{y} T ⟶^{b} B) ↦ (X × Y ⟵^{x×y} S × T ⟶^{a×b} A × B) ,

and its nullary aspect provided by

1 ⟵^{1} 1 ⟶^{1} 1 ,

The authors gratefully acknowledge financial support from the Australian ARC, the Italian CNR and the Canadian NSERC. Diagrams typeset using M. Barr’s diagram package, diagxy.tex.

Received by the editors 2009-10-19 and, in revised form, 2010-02-10.

Transmitted by Robert Paré. Published on 2010-02-11.

2000 Mathematics Subject Classification: 18A25.

Key words and phrases: bicategory, finite products, discrete, comonad, Eilenberg-Moore object.

\begin{matrix}c \\ ◯ \end{matrix} Stephen Lack, R.F.C Walters, and R.J. Wood, 2010. Permission to copy for private use granted.

the terminal object of Span E(1, 1). Both of these lax functors are readily seen to be pseudofunctors so that Span E is a cartesian bicategory as in [CKWW].

The purpose of this paper is to characterize those cartesian bicategories B which are biequivalent to Span E, for some category E with finite limits. Certain aspects of a solution to the problem are immediate. A biequivalence B ∼ Span E provides

MapB ∼ MapSpan E ∼ E

so that we must ensure firstly that MapB is essentially locally discrete. From the characterization of bicategories of relations as locally ordered cartesian bicategories in [C&W] one suspects that the following axiom will figure prominently in providing essential local discreteness for MapB.

1.1. Axiom. Frobenius: A cartesian bicategory B is said to satisfy the Frobenius axiom if, for each A in B, A is Frobenius. Indeed Frobenius objects in cartesian bicategories were defined and studied in [W&W] where amongst other things it is shown that if A is Frobenius in cartesian B then, for all X, MapB(X, A) is a groupoid. (This theorem was generalized considerably in [LSW] which explained further aspects of the Frobenius concept.) However, essential local discreteness for MapB requires also that the MapB(X, A) be ordered sets (which is automatic for locally ordered B). Here we study also separable objects in cartesian bicategories for which we are able to show that if A is separable in cartesian B then, for all X, MapB(X, A) is an ordered set and a candidate axiom is:

1.2. Axiom. Separable: A cartesian bicategory B is said to satisfy the Separable axiom if, for each A in B, A is separable.

In addition to essential local discreteness, it is clear that we will need an axiom which provides tabulation of each arrow of B by a span of maps. Since existence of Eilenberg-Moore objects is a basic 2-dimensional limit concept, we will express tabulation in terms of this requirement; we note that existence of pullbacks in MapB follows easily from tabulation. In the bicategory Span E, the comonads G:A [figure] A are precisely the symmetric spans g : A [figure] X [figure] A : g; the map g : X [figure] A together with gη_{g} : g [figure] gg^{∗}g provides an Eilenberg-Moore coalgebra for g:A [figure] X [figure] A:g. We will posit:

1.3. Axiom. Eilenberg-Moore for Comonads: Each comonad (A, G) in B has an Eilenberg-Moore object.

Conversely, any map (left adjoint) g: X [figure] A in Span E provides an Eilenberg-Moore object for the comonad gg^{∗}. We further posit:

1.4. Axiom. Maps are Comonadic: Each left adjoint g:X [figure] A in B is comonadic. from which, in our context, we can also deduce the Frobenius and Separable axioms.

In fact we shall also give, in Proposition 3.1 below, a straightforward proof that MapB is locally essentially discrete whenever Axiom 1.4 holds. But due to the importance of the Frobenius and separability conditions in other contexts, we have chosen to analyze them in their own right.

2. Preliminaries

We recall from [CKWW] that a bicategory B (always, for convenience, assumed to be normal) is said to be cartesian if the subbicategory of maps (by which we mean left adjoint arrows), M = MapB, has finite products − × − and 1; each hom-category B(B, C) has finite products − ∧ − and ⊤; and a certain derived tensor product − ⊗ − and I on B, extending the product structure of M, is functorial. As in [CKWW], we write p and r for the first and second projections at the global level, and similarly π and ρ for the projections at the local level. If f is a map of B — an arrow of M — we will write η_{f}, ϵ_{f}:f ⊣ f^{∗} for a chosen adjunction in B that makes it so. It was shown that the derived tensor product of a cartesian bicategory underlies a symmetric monoidal bicategory structure. We recall too that in [W&W] Frobenius objects in a general cartesian bicategory were defined and studied. We will need the central results of that paper too. Throughout this paper, B is assumed to be a cartesian bicategory.

As in [CKWW] we write

[figure: M; M; G = GroB; M; ∂_{0}; G = GroB; M; ∂_{1}]

for the Grothendieck span corresponding to

M^{op} × M [figure: i^{op}×i] B^{op} × B [figure: B(−,−)] CAT

where i:M [figure] B is the inclusion. A typical arrow of G, (f, α, u):(X, R, A) [figure] (Y, S, B) can be depicted by a square

[figure: A; B; u; X; A; R; X; Y; f; Y; B; S; α]  (1)

and such arrows are composed by pasting. A 2-cell (ϕ, ψ):(f, α, u) [figure] (g, β, v) in G is a pair of 2-cells ϕ: f [figure] g, ψ: u [figure] v in M which satisfy the obvious equation. The (strict) pseudofunctors ∂_{0} and ∂_{1} should be regarded as domain and codomain respectively. Thus, applied to (1), ∂_{0} gives f and ∂_{1} gives u. The bicategory G also has finite products, which are given on objects by − ⊗ − and I; these are preserved by ∂_{0} and ∂_{1}.

The Grothendieck span can also be thought of as giving a double category (of a suitably weak flavour), although we shall not emphasize that point of view.

2.1. The arrows of G are particularly well suited to relating the various product structures in a cartesian bicategory. In 3.31 of [CKWW] it was shown that the local binary product, for R, S:X [figure] A, can be recovered to within isomorphism from the defined tensor product by

R ∧ S ≅ d_{A}^{∗}(R ⊗ S)d_{X}

A slightly more precise version of this is that the mate of the isomorphism above, with respect to the single adjunction d_{A} ⊣ d_{A}^{∗}, defines an arrow in G

[figure: A; A ⊗ A; d_{A}; X; A; R∧S; X; X ⊗ X; d_{X}; X ⊗ X; A ⊗ A; R⊗S]

which when composed with the projections of G, recovers the local projections as in

[figure: A; A ⊗ A; d_{A}; X; A; R∧S; X; X ⊗ X; d_{X}; X ⊗ X; A ⊗ A; R⊗S; A ⊗ A; A; p_{A,A}; X ⊗ X; A ⊗ A; R⊗S; X ⊗ X; X; p_{X,X}; X; A; R; p̃_{R,S}; ≅; A; A; 1_{A}; X; A; R∧S; X; X; 1_{X}; X; A; R; π]

for the first projection, and similarly for the second. The unspecified ≅ in G is given by a pair of convenient isomorphisms p_{X,X}d_{X} ≅ 1_{X} and p_{A,A}d_{A} ≅ 1_{A} in M. Similarly, when R ∧ S [figure] R ⊗ S is composed with (r_{X,X}, r̃_{R,S}, r_{A,A}) the result is (1_{X}, ρ, 1_{A}):R ∧ S [figure] S.

2.2. Quite generally, an arrow of G as given by the square (1) will be called a commutative square if α is invertible. An arrow of G will be said to satisfy the Beck condition if the mate of α under the adjunctions f ⊣ f^{∗} and u ⊣ u^{∗}, as given in the square below (no longer an arrow of G), is invertible.

[figure: A; B; u^{∗}; X; A; R; X; Y; f^{∗}; Y; B; S; α^{∗}]

Thus Proposition 4.7 of [CKWW] says that projection squares of the form p̃_{R,1_{Y}} and r̃_{1_{X},S} are commutative while Proposition 4.8 of [CKWW] says that these same squares satisfy the Beck condition. If R and S are also maps and α is invertible then α^{−1} gives rise to another arrow of G, from f to u with reference to the square above, which may or may not satisfy the Beck condition. The point here is that a commutative square of maps gives rise to two, generally distinct, Beck conditions. It is well known that, for bicategories of the form Span E and Rel E, all pullback squares of maps satisfy both Beck conditions. A category with finite products has automatically a number of pullbacks which we might call product-absolute pullbacks because they are preserved by all functors which preserve products. In [W&W] the Beck conditions for the product-absolute pullback squares of the form

[figure: A; A × A; d; A × A; A; d; A × A; A × A × A; d×A; A × A × A; A × A; A×d]

were investigated. (In fact, in this case it was shown that either Beck condition implies the other.) The objects for which these conditions are met are called Frobenius objects.

2.3. Proposition. For a cartesian bicategory, the axiom Maps are Comonadic implies the axiom Frobenius.

Proof. It suffices to show that the 2-cell δ_{1} below is invertible:

[figure: A ⊗ A; A ⊗ (A ⊗ A); 1⊗d^{∗}; A; A ⊗ A; d; A; A ⊗ A; d^{∗}; A ⊗ A; A ⊗ (A ⊗ A); A ⊗ A; (A ⊗ A) ⊗ A; d⊗1; (A ⊗ A) ⊗ A; A ⊗ (A ⊗ A); a; δ_{1}; A; A ⊗ A; d^{∗}; A ⊗ A; A; r; A ⊗ A; A ⊗ (A ⊗ A); 1⊗d^{∗}; A ⊗ (A ⊗ A); A ⊗ A; r; r̃_{1_{A},d^{∗}}]

The paste composite of the squares is invertible (being essentially the identity 2-cell on d^{∗}). The lower 2-cell is invertible by Proposition 4.7 of [CKWW] so that the whisker composite rδ_{1} is invertible. Since r is a map it reflects isomorphisms, by Maps are Comonadic, and hence δ_{1} is invertible.

2.4. Remark. It was shown in [W&W] that, in a cartesian bicategory, the Frobenius objects are closed under finite products. It follows that the full subbicategory of a cartesian bicategory determined by the Frobenius objects is a cartesian bicategory which satisfies the Frobenius axiom.

3. Separable Objects and Discrete Objects in Cartesian Bicategories

In this section we look at separability for objects of cartesian bicategories. Since for an object A which is both separable and Frobenius, the hom-category MapB(X, A) is essentially discrete, for all X, we shall then be able to show that MapB is essentially discrete by showing that all objects in B are separable and Frobenius. But first we record the following direct argument:

3.1. Proposition. If B is a bicategory in which all maps are comonadic and MapB has a terminal object, then MapB is locally essentially discrete.

Proof. We must show that for all objects X and A, the hom-category MapB(X, A) is essentially discrete. As usual, we write 1 for the terminal object of MapB and t_{A} : A [figure] 1 for the essentially unique map, which by assumption is comonadic. Let f, g : X [figure] A be maps from X to A. If α : f [figure] g is any 2-cell, then t_{A}α is invertible, since 1 is terminal in MapB. But since t_{A} is comonadic, it reflects isomorphisms, and so α is invertible. Furthermore, if β : f [figure] g is another 2-cell, then t_{A}α = t_{A}β by the universal property of 1 once again, and now α = β since t_{A} is faithful. Thus there is at most one 2-cell from f to g, and any such 2-cell is invertible.

In any (bi)category with finite products the diagonal arrows d_{A}:A [figure] A × A are (split) monomorphisms so that in the bicategory M the following square is a product-absolute pullback

[figure: A; A ⊗ A; d_{A}; A; A; 1_{A}; A; A; 1_{A}; A; A ⊗ A; d_{A}]

that gives rise to a single G arrow.

3.2. Definition. An object A in a cartesian bicategory is said to be separable if the G arrow above satisfies the Beck condition.

Of course the invertible mate condition here says precisely that the unit η_{d_{A}}:1_{A} [figure] d_{A}^{∗}d_{A} for the adjunction d_{A} ⊣ d_{A}^{∗} is invertible. Thus Axiom 1.2, as stated in the Introduction, says that, for all A in B, η_{d_{A}} is invertible.

3.3. Remark. For a map f it makes sense to define f is fully faithful to mean that η_{f} is invertible. For a category A the diagonal d_{A} is fully faithful if and only if A is an ordered set.

3.4. Proposition. For an object A in a cartesian bicategory, the following are equivalent:

i) A is separable;

ii) for all f:X [figure] A in M, the diagram f [figure] f [figure] f is a product in B(X, A);

iii) 1_{A} [figure] 1_{A} [figure] 1_{A} is a product in B(A, A);

iv) 1_{A} [figure] ⊤_{A,A} is a monomorphism in B(A, A);

v) for all G [figure] 1_{A} in B(A, A), the diagram G [figure] G [figure] 1_{A} is a product in B(A, A).

Proof. [i) ⟹ ii)] A local product of maps is not generally a map but here we have:

f ∧ f ≅ d_{A}^{∗}(f ⊗ f)d_{X} ≅ d_{A}^{∗}(f × f)d_{X} ≅ d_{A}^{∗}d_{A}f ≅ f

[ii) ⟹ iii)] is trivial.

[iii) ⟹ i)] Note the use of pseudo-functoriality of ⊗:

d_{A}^{∗}d_{A} ≅ d_{A}^{∗}1_{A⊗A}d_{A} ≅ d_{A}^{∗}(1_{A} ⊗ 1_{A})d_{A} ≅ 1_{A} ∧ 1_{A} ≅ 1_{A}

[iii) ⟹ iv)] To say that 1_{A} [figure] 1_{A} [figure] 1_{A} is a product in B(A, A) is precisely to say that

[figure: 1_{A}; ⊤_{A,A}; 1_{A}; 1_{A}; 1_{1_{A}}; 1_{A}; 1_{A}; 1_{1_{A}}; 1_{A}; ⊤_{A,A}]

is a pullback in B(A, A) which in turn is precisely to say that 1_{A} [figure] ⊤_{A,A} is a monomorphism in B(A, A)

[iv) ⟹ v)] It is a generality that if an object S in a category is subterminal then for any G [figure] S, necessarily unique, G [figure] G [figure] S is a product diagram.

[v) ⟹ iii)] is trivial.

3.5. Corollary. [Of iv)] For a cartesian bicategory, the axiom Maps are Comonadic implies the axiom Separable.

Proof. We have ⊤_{A,A} = t_{A}^{∗}t_{A} for the map t_{A}:A [figure] 1. It follows that the unique 1_{A} [figure] t_{A}^{∗}t_{A} is η_{t_{A}}. Since t_{A} is comonadic, η_{t_{A}} is the equalizer shown:

1_{A} [figure: η_{t_{A}}] t_{A}^{∗}t_{A} [figure: t_{A}^{∗}t_{A}η_{t_{A}}; η_{t_{A}}t_{A}^{∗}t_{A}] t_{A}^{∗}t_{A}t_{A}^{∗}t_{A}

and hence a monomorphism.

3.6. Corollary. [Of iv)] For separable A in cartesian B, an arrow G:A [figure] A admits at most one copoint G [figure] 1_{A} depending upon whether the unique arrow G [figure] ⊤_{A,A} factors through 1_{A} [figure] ⊤_{A,A}.

3.7. Proposition. In a cartesian bicategory, the separable objects are closed under finite products.

Proof. If A and B are separable objects then applying the homomorphism ⊗:B×B [figure] B we have an adjunction d_{A} × d_{B} ⊣ d_{A}^{∗} ⊗ d_{B}^{∗} with unit η_{d_{A}} ⊗ η_{d_{B}} which being an isomorph of the adjunction d_{A⊗B} ⊣ d_{A⊗B}^{∗} with unit η_{d_{A⊗B}} (via middle-four interchange) shows that the separable objects are closed under binary products. On the other hand, d_{I} is an equivalence so that I is also separable.

3.8. Corollary. For a cartesian bicategory, the full subbicategory determined by the separable objects is a cartesian bicategory which satisfies the axiom Separable.

3.9. Proposition. If A is a separable object in a cartesian bicategory B, then, for all X in B, the hom-category M(X, A) is an ordered set, meaning that the category structure forms a reflexive, transitive relation.

Proof. Suppose that we have arrows α, β:g [figure] f in M(X, A). In B(X, A) we have

[figure: f; f; g; f; α; g; f; β; g; f ∧ f; γ; f ∧ f; f; π; f ∧ f; f; ρ]

By Proposition 3.4 we can take f ∧ f = f and π = 1_{f} = ρ so that we have α = γ = β. It follows that M(X, A) is an ordered set.

3.10. Definition. An object A in a cartesian bicategory is said to be discrete if it is both Frobenius and separable. We write DisB for the full subbicategory of B determined by the discrete objects.

3.11. Remark. Beware that this is quite different to the notion of discreteness in a bicategory. An object A of a bicategory is discrete if each hom-category B(X, A) is discrete; A is essentially discrete if each B(X, A) is equivalent to a discrete category. The notion of discreteness for cartesian bicategories defined above turns out to mean that A is essentially discrete in the bicategory MapB.

From Proposition 3.7 above and Proposition 3.4 of [W&W] we immediately have

3.12. Proposition. For a cartesian bicategory B, the full subbicategory DisB of discrete objects is a cartesian bicategory in which every object is discrete.

And from Proposition 3.9 above and Theorem 3.13 of [W&W] we have

3.13. Proposition. If A is a discrete object in a cartesian bicategory B then, for all X in B, the hom category M(X, A) is an equivalence relation.

If both the Frobenius axiom of [W&W] and the Separable axiom of this paper hold for our cartesian bicategory B, then every object of B is discrete. In this case, because M is a bicategory, the equivalence relations M(X, A) are stable under composition from both sides. Thus writing |M(X, A)| for the set of objects of M(X, A) we have a mere category, E whose objects are those of M (and hence also those of B) and whose hom sets are the quotients |M(X, A)|/M(X, A). If the E(X, A) are regarded as discrete categories, so that E is a locally discrete bicategory then the functors M(X, A) [figure] |M(X, A)|/M(X, A) constitute the effect on homs functors for an identity on objects biequivalence M [figure] E. To summarize

3.14. Theorem. If a cartesian bicategory B satisfies both the Frobenius and Separable axioms then the bicategory of maps M is biequivalent to the locally discrete bicategory E.

In the following lemma we show that any copointed endomorphism of a discrete object can be made into a comonad; later on, we shall see that this comonad structure is in fact unique.

3.15. Lemma. If A is a discrete object in a cartesian bicategory B then, for any copointed endomorphism arrow ϵ:G [figure] 1_{A}:A [figure] A, there is a 2-cell δ = δ_{G}:G [figure] GG satisfying

[figure: G; G; G; G; 1; G; G; 1; G; GG; δ; GG; G; Gϵ; GG; G; ϵG]

and if both G, H:A [figure] A are copointed, so that GH:A [figure] A is also copointed, and ϕ:G [figure] H is any 2-cell, then the δ’s satisfy

[figure: GG; HH; ϕϕ; G; GG; δ; G; H; ϕ; H; HH; δ; and; GH; GH; 1; GHGH; GH; δ; GHGH; GH; (Gϵ)(ϵH)]

Proof. We define δ = δ_{G} to be the pasting composite

[figure: A; AA; d; A; AAA; d_{3}; AA; AAA; 1d; AAA; AAA; G11; AA; AAA; 1d; AA; AA; G1; AA; AAA; 1d; AAA; AA; d^{∗}1; AA; AAA; 1d; AA; A; d^{∗}; A; AA; d; AAA; AAA; G11; AAA; AAA; GGG; AAA; AAA; 11G; AAA; AA; d^{∗}1; AAA; AAA; 11G; AAA; AA; d^{∗}1; AA; AA; 1G; AAA; AA; d^{∗}1; AAA; A; d_{3}^{∗}; AA; A; d^{∗}; A; A; G∧G∧G; A; A; G; A; A; G; A; A; G; GϵG; δ_{3}; 1; 2; 3; 4; 5; 6]

wherein ⊗ has been abbreviated by juxtaposition and all subregions not explicitly inhabited by a 2-cell are deemed to be inhabited by the obvious invertible 2-cell. A reference number has been assigned to those invertible 2-cells which arise from the hypotheses. As in [W&W], d_{3}’s denote 3-fold diagonal maps and, similarly, we write δ_{3} for a local 3-fold diagonal.

The invertible 2-cell labelled by ‘1’ is that defining A to be Frobenius. The 3-fold composite of arrows in the region labelled by ‘2’ is G ∧ 1_{A} and, similarly, in that labelled by ‘3’ we have 1_{A} ∧ G. Each of these is isomorphic to G because A is separable and G is copointed. The isomorphisms in ‘4’ and ‘5’ express the pseudo-functoriality of ⊗ in the cartesian bicategory B. Finally ‘6’ expresses the ternary local product in terms of the ternary ⊗ as in [W&W]. Demonstration of the equations is effected easily by pasting composition calculations.

3.16. Theorem. If G and H are copointed endomorphisms on a discrete A in a cartesian B then

G [figure: Gϵ] GH [figure: ϵH] H

is a product diagram in B(A, A).

Proof. If we are given α:K [figure] G and β:K [figure] H then K is also copointed and we have

K [figure: δ] KK [figure: αβ] GH

as a candidate pairing. That this candidate satisfies the universal property follows from the equations of Lemma 3.15 which are precisely those in the equational description of binary products. We remark that the ‘naturality’ equations for the projections follow immediately from uniqueness of copoints.

3.17. Corollary. If A is discrete in a cartesian B, then an endo-arrow G: A [figure] A admits a comonad structure if and only if G has the copointed property, and any such comonad structure is unique.

Proof. The Theorem shows that the arrow δ: G [figure] GG constructed in Lemma 3.15 is the product diagonal on G in the category B(A, A) and, given ϵ:G [figure] 1_{A}, this is the only comonad comultiplication on G.

3.18. Remark. It is clear that 1_{A} is terminal with respect to the copointed objects in B(A, A).

3.19. Proposition. If an object B in a bicategory B has 1_{B} subterminal in B(B, B) then, for any map f:A [figure] B, f is subterminal in B(A, B) and f^{∗} is subterminal in B(B, A). In particular, in a cartesian bicategory in which every object is separable, every adjoint arrow is subterminal.

Proof. Precomposition with a map preserves terminal objects and monomorphisms, as does postcomposition with a right adjoint.

4. Bicategories of Comonads

The starting point of this section is the observation, made in the introduction, that a comonad in the bicategory Span E is precisely a span of the form

A ⟵^{g} X ⟶^{g} A

in which both legs are equal.

We will write C = ComB for the bicategory of comonads in B, Com being one of the duals of Street’s construction Mnd in [ST]. Thus C has objects given by the comonads (A, G) of B. The structure 2-cells for comonads will be denoted ϵ = ϵ_{G}, for counit and δ = δ_{G}, for comultiplication. An arrow in C from (A, G) to (B, H) is a pair (F, ϕ) as shown in

[figure: A; B; F; A; A; G; A; B; F; B; B; H; ϕ]

satisfying

[figure: F 1_{A}; 1_{B}F; =; F G; F 1_{A}; F ϵ; F G; HF; ϕ; HF; 1_{B}F; ϵF; and; F GG; HF G; ϕG; F G; F GG; F δ; F G; HF G; HF G; HHF; Hϕ; HF G; HF; HF; HHF; δF; F G; HF; ϕ]  (2)

(where, as often, we have suppressed the associativity constraints of our normal, cartesian, bicategory B). A 2-cell τ:(F, ϕ) [figure] (F′, ϕ′):(A, G) [figure] (B, H) in C is a 2-cell τ:F [figure] F′ in B satisfying

[figure: F′G; HF′; ϕ′; F G; F′G; τG; F G; HF; ϕ; HF; HF′; Hτ]  (3)

There is a pseudofunctor I:B [figure] C given by

I(τ:F [figure] F′:A [figure] B) = τ:(F, 1_{F}) [figure] (F′, 1_{F′}):(A, 1_{A}) [figure] (B, 1_{B})

From [ST] it is well known that a bicategory B has Eilenberg-Moore objects for comonads if and only if I:B [figure] C has a right biadjoint, which we will denote by E:C [figure] B. We write E(A, G) = A_{G} and the counit for I ⊣ E is denoted by

[figure: A_{G}; A; g_{G}; A_{G}; A_{G}; 1_{A_{G}}; A_{G}; A; g_{G}; A; A; G; γ_{G}; A_{G}; A; g_{G}; A; A_{G}; g_{G}; A; A; G; or, using normality of B, better by; γ_{G}]

with (g_{G}, γ_{G}) abbreviated to (g, γ) when there is no danger of confusion. It is standard that each g = g_{G} is necessarily a map (whence our lower case notation) and the mate gg^{∗} [figure] G of γ is an isomorphism which identifies ϵ_{g} and ϵ_{G}.

We will write D for the locally full subbicategory of C determined by all the objects and those arrows of the form (f, ϕ), where f is a map, and write j:D [figure] C for the inclusion. It is clear that the pseudofunctor I:B [figure] C restricts to give a pseudofunctor J:M [figure] D. We say that the bicategory B has Eilenberg-Moore objects for comonads, as seen by M, if J:M [figure] D has a right biadjoint. (In general, this property does not follow from that of Eilenberg-Moore objects for comonads.)

4.1. Remark. In the case B = Span E, a comonad in B can, as we have seen, be identified with a morphism in E. This can be made into the object part of a biequivalence between the bicategory D and the category E^{2} of arrows in E. If we further identify M with E, then the inclusion j : D [figure] C becomes the diagonal E [figure] E^{2}; of course this does have a right adjoint, given by the domain functor.

4.2. Theorem. If B is a cartesian bicategory in which every object is discrete, the bicategory D = D(B) admits the following simpler description:

i) An object is a pair (A, G) where A is an object of B and G:A [figure] A admits a copoint;

ii) An arrow (f, ϕ):(A, G) [figure] (B, H) is a map f:A [figure] B and a 2-cell ϕ:fG [figure] Hf;

iii) A 2-cell τ:(f, ϕ) [figure] (f′, ϕ′):(A, G) [figure] (B, H) is a 2-cell satisfying τ:f [figure] f′ satisfying equation (3).

Proof. We have i) by Corollary 3.17 while iii) is precisely the description of a 2-cell in D, modulo the description of the domain and codomain arrows. So, we have only to show ii), which is to show that the equations (2) hold automatically under the hypotheses. For the first equation of (2) we have uniqueness of any 2-cell fG [figure] f because f is subterminal by Proposition 3.19. For the second, observe that the terminating vertex, HHf, is the product Hf ∧ Hf in M(A, B) because HH is the product H ∧ H in M(B, B) by Theorem 3.16 and precomposition with a map preserves all limits. For HHf seen as a product, the projections are, again by Theorem 3.16, Hϵf and ϵHf. Thus, it suffices to show that the diagram for the second equation commutes when composed with both Hϵf and ϵHf. We have

[figure: fGG; HfG; ϕG; fG; fGG; fδ; fG; HfG; HfG; HHf; Hϕ; HfG; Hf; Hf; HHf; δf; fG; Hf; ϕ; HfG; HHf; Hϕ; HfG; Hf; Hfϵ; HHf; Hf; Hϵf; fG; Hf; ϕ; fGG; fG; fGϵ; fGG; HfG; ϕG; fG; fGG; fδ; fG; HfG; HfG; HHf; Hϕ; HfG; Hf; Hf; HHf; δf; fG; Hf; ϕ; fGG; HfG; ϕG; fGG; fG; fϵG; HfG; fG; ϵfG; fG; Hf; ϕ; HHf; Hf; ϵHf]

in which each of the lower triangles commutes by the first equation of (2) already established. Using comonad equations for G and H, it is obvious that each composite is ϕ.

Finally, let us note that D is a subbicategory, neither full nor locally full, of the Grothendieck bicategory G and write K:D [figure] G for the inclusion. We also write ι:M [figure] G for the composite pseudofunctor KJ. Summarizing, we have introduced the following commutative diagram of bicategories and pseudofunctors

[figure: B; C; I; M; B; i; M; D; J; D; C; j; D; G; K; M; G; ι] ;

note also that in our main case of interest B = Span E, each of M, D, and G is biequivalent to a mere category. Ultimately, we are interested in having a right biadjoint, say τ, of ι. For such a biadjunction ι ⊣ τ the counit at an object R:X [figure] A in G will take the form

[figure: τR; A; v_{R}; X; τR; u_{R}; X; A; R; ω_{R}]  (4)

(where, as for a biadjunction I ⊣ E:C [figure] B, a triangle rather than a square can be taken as the boundary of the 2-cell by the normality of B). In fact, we are interested in the case where we have ι ⊣ τ and moreover the counit components ω_{R}:v_{R} [figure] Ru_{R} enjoy the property that their mates v_{R}u_{R}^{∗} [figure] R with respect to the adjunction u_{R} ⊣ u_{R}^{∗} are invertible. In this way we represent a general arrow of B in terms of a span of maps. Since biadjunctions compose we will consider adjunctions J ⊣ F and K ⊣ G and we begin with the second of these.

4.3. Theorem. For a cartesian bicategory B in which every object is discrete, there is an adjunction K ⊣ G: G [figure] D where, for R: X [figure] A in G, the comonad G(R) and its witnessing copoint ϵ: G(R) [figure] 1_{XA} are given by the left diagram below and the counit μ:KG(R) [figure] R is given by the right diagram below, all in notation suppressing ⊗:

[figure: XA; XA; 1_{XA}; XA; XXA; dA; XA; XXA; p_{1,3}; ≃; XXA; XAA; XRA; XA; XA; 1_{XA}; p̃_{1,3}; XA; XA; 1_{XA}; XAA; XA; Xd^{∗}; XAA; XA; p_{1,3}; XA; XA; G(R); XA; X; p; XA; XXA; dA; X; XXA; p_{2}; XXA; XAA; XRA; X; A; R; XA; A; r; XAA; XA; Xd^{∗}; XAA; A; p_{2}; ≃; p̃_{2}; XA; XA; G(R)]

Moreover, the mate rG(R)p^{∗} [figure] R of the counit μ is invertible.

In the left diagram, the p_{1,3} collectively denote projection from the three-fold product in G to the product of the first and third factors. In the right diagram, the p_{2} collectively denote projection from the three-fold product in G to the second factor. The upper triangles of the two diagrams are the canonical isomorphisms. The lower left triangle is the mate of the canonical isomorphism 1 [figure: ≃] p_{1,3}(Xd). The lower right triangle is the mate of the canonical isomorphism r [figure: ≃] p_{2}(Xd).

Proof. Given a comonad H:T [figure] T and an arrow

[figure: T; A; a; T; T; H; T; X; x; X; A; R; ψ]

in G, we verify the adjunction claim by showing that there is a unique arrow

[figure: T; XA; f; T; T; H; T; XA; f; XA; XA; G(R); ϕ]

in D, whose composite with the putative counit μ is (x, ψ, a). It is immediately clear that the unique solution for f is (x, a) and to give ϕ:(x, a)H [figure] Xd^{∗}(XRA)dA(x, a) is to give the mate Xd(x, a)H [figure] (XRA)dA(x, a) which is (x, a, a)H [figure] (XRA)(x, x, a) and can be seen as a G arrow:

[figure: T; XAA; (x,a,a); T; T; H; T; XXA; (x,x,a); XXA; XAA; XRA; (α,β,γ)]

where we exploit the description of products in G. From this description it is clear, since p̃_{2}(α, β, γ) = β as a composite in G, that the unique solution for β is ψ. We have seen in Theorem 3.17 that the conditions (2) hold automatically in D under the assumptions of the Theorem. From the first of these we have:

[figure: T; XAA; (x,a,a); T; T; H; T; XXA; (x,x,a); XXA; XAA; XRA; (α,β,γ); XAA; XA; p_{1,3}; XXA; XAA; XRA; XXA; XA; p_{1,3}; XA; XA; 1_{XA}; p_{1,3}; =; T; XA; (x,a); T; T; 1_{T}; T; XA; (x,a); XA; XA; 1_{XA}; κ_{(x,a)}; T; T; H; ϵ_{H}]

So, with a mild abuse of notation, we have (α, γ) = (1_{x}ϵ_{H}, 1_{a}ϵ_{H}), uniquely, and thus the unique solutions for α and γ are 1_{x}ϵ_{H} and 1_{a}ϵ_{H} respectively. This shows that ϕ is necessarily the mate under the adjunctions considered of (1_{x}ϵ_{H}, ψ, 1_{a}ϵ_{H},). Since D and G are essentially locally discrete this suffices to complete the claim that K ⊣ G. It only remains to show that the mate rG(R)p^{∗} [figure] R of the counit μ is invertible. In the three middle squares of the diagram

[figure: XA; XA; G(R); XXA; XX; p^{∗}; XA; XXA; dA; XA; X; p^{∗}; X; XX; d; p̃_{d,1_{A}}^{∗}; XAA; XA; p^{∗}; XXA; XAA; XRA; XXA; XX; p^{∗}; XX; XA; XR; p̃_{XR,1_{A}}^{∗}; XA; A; r; XAA; XA; Xd^{∗}; XAA; XA; p^{∗}; XA; A; r; ≃; XA; A; r; XX; XA; XR; XX; X; r; X; A; R; r_{1_{X} ,R}; X; X; 1_{X}; A; A; 1_{A}]

the top two are invertible 2-cells by Proposition 4.18 of [CKWW] while the lower one is the obvious invertible 2-cell constructed from Xd^{∗}p^{∗} ≅ 1_{X,A}. The right square is an invertible 2-cell by Proposition 4.17 of [CKWW]. This shows that the mate rG(R)p^{∗} [figure] R of μ is invertible.

4.4. Remark. It now follows that the unit of the adjunction K ⊣ G is given (in notation suppressing ⊗) by:

[figure: T; T T; d; T; T T T; d_{3}; T T; T T T; dT; T; T T; d; T T T; T; d_{3}; T T T; T T; T d^{∗}; T; T; H; T T T; T T T; HHH; T T T; T T T; T HT; ≃; d̃_{3}; ϵHϵ; ≃]

where the d_{3} collectively denote 3-fold diagonalization (1, 1, 1) in G. The top triangle is a canonical isomorphism while the lower triangle is the mate of the canonical isomorphism (T ⊗ d)d [figure: ≃] d_{3} and is itself invertible, by separability of T.

Before turning to the question of an adjunction J ⊣ F, we note:

4.5. Lemma. In a cartesian bicategory in which Maps are Comonadic, if gF ≅ h with g and h maps, then F is also a map.

Proof. By Theorem 3.11 of [W&W] it suffices to show that F is a comonoid homomorphism, which is to show that the canonical 2-cells t̃_{F}:tF [figure] t and d̃_{F}:dF [figure] (F ⊗ F)d are invertible. For the first we have:

tF ≅ tgF ≅ th ≅ t

Simple diagrams show that we do get the right isomorphism in this case and also for the next:

(g ⊗ g)(dF) ≅ dgF ≅ dh ≅ (h ⊗ h)d ≅ (g ⊗ g)(F ⊗ F)d

which gives dF ≅ (F ⊗ F)d since the map g ⊗ g reflects isomorphisms.

4.6. Theorem. If B is a cartesian bicategory which has Eilenberg-Moore objects for Comonads and for which Maps are Comonadic then B has Eilenberg-Moore objects for Comonads as Seen by M, which is to say that J:M [figure] D has a right adjoint. Moreover, the counit for the adjunction, say JF [figure] 1_{D}, necessarily having components of the form γ:g [figure] Gg with g a map, has gg^{∗} [figure] G invertible.

Proof. It suffices to show that the adjunction I ⊣ E:C [figure] B restricts to J ⊣ F:D [figure] M. For this it suffices to show that, given (h, θ):JT [figure] (A, G), the F :T [figure] A_{G} with gF ≅ h which can be found using I ⊣ E has F a map. This follows from Lemma 4.5.

4.7. Theorem. A cartesian bicategory which has Eilenberg-Moore objects for Comonads and for which Maps are Comonadic has tabulation in the sense that the inclusion ι:M [figure] G has a right adjoint τ and the counit components ω_{R}:v_{R} [figure] Ru_{R} as in (4) have the property that the mates v_{R}u_{R}^{∗} [figure] R, with respect to the adjunctions u_{R} ⊣ u_{R}^{∗}, are invertible.

Proof. Using Theorems 4.3 and 4.6 we can construct the adjunction ι ⊣ τ by composing J ⊣ F with K ⊣ G. Moreover, the counit for ι ⊣ τ is the pasting composite:

[figure: T; X ⊗ A; (u,v); X ⊗ A; T; (u,v); X ⊗ A; X ⊗ A; G(R); T; X; u; T; A; v; γ; X ⊗ A; A; r; X ⊗ A; X ⊗ A; G(R); X ⊗ A; X; p; X; A; R; μ]

where the square is the counit for K ⊣ G; and the triangle, the counit for J ⊣ F, is an Eilenberg-Moore coalgebra for the comonad G(R). The arrow component of the Eilenberg-Moore coalgebra is necessarily of the form (u, v), where u and v are maps, and it also follows that we have (u, v)(u, v)^{∗} ≅ G(R). Thus we have

vu^{∗} ≅ r(u, v)(p(u, v))^{∗} ≅ r(u, v)(u, v)^{∗}p^{∗} ≅ rG(R)p^{∗} ≅ R

where the first two isomorphisms are trivial, the third arises from the invertibility of the mate of γ as an Eilenberg-Moore structure, and the fourth is invertibility of μ, as in Theorem 4.3.

4.8. Theorem. For a cartesian bicategory B with Eilenberg-Moore objects for Comonads and for which Maps are Comonadic, MapB has pullbacks satisfying the Beck condition (meaning that for a pullback square

[figure: N; A; a; P; N; p; P; M; r; M; A; b; ≃]  (5)

the mate pr^{∗} [figure] a^{∗}b of ap ≅ br in B, with respect to the adjunctions r ⊣ r^{∗} and a ⊣ a^{∗}, is invertible).

Proof. Given the cospan a:N [figure] A [figure] M:b in MapB, let P together with (r, σ, p) be a tabulation for a^{∗}b:M [figure] N. Then pr^{∗} [figure] a^{∗}b, the mate of σ:p [figure] a^{∗}br with respect to r ⊣ r^{∗}, is invertible by Theorem 4.7. We have also ap [figure] br, the mate of σ:p [figure] a^{∗}br with respect to a ⊣ a^{∗}. Since A is discrete, ap [figure] br is also invertible and is the only 2-cell between the composite maps in question. If we have also u: N [figure] T [figure] M: v, for maps u and v with au ≅ bv, then the mate u [figure] a^{∗}bv ensures that the span u:N [figure] T [figure] M:v factors through P by an essentially unique map w:T [figure] P with pw ≅ u and rw ≅ v.

4.9. Proposition. In a cartesian bicategory with Eilenberg-Moore objects for Comonads and for which Maps are Comonadic, every span of maps x:X [figure] S [figure] A:a gives rise to the following tabulation diagram:

[figure: S; A; a; X; S; x; X; A; ax^{∗}; aη_{x}]

Proof. A general tabulation counit ω_{R} : v_{R} [figure] Ru_{R} is given in terms of the Eilenberg-Moore coalgebra for the comonad (u, v)(u, v)^{∗} and necessarily (u, v)(u, v)^{∗} ≅ G(R). It follows that for R = ax^{∗}, it suffices to show that G(ax^{∗}) ≅ (x, a)(x, a)^{∗}. Consider the diagram (with ⊗ suppressed):

[figure: XXA; XAA; XSA; XXA; XxA; XSA; XAA; XaA; SA; XSA; (x,S)A; SA; XS; XS; XSA; X(S,a); SA; XS; S; SA; (S,a); S; XS; (x,S); XA; XXA; dA; SA; XA; xA; SA; XXA; XA; XAA; Xd; XS; XA; Xa; XS; XAA; S; XA; (x,a); S; XA; (x,a)]

The comonoid G(ax^{∗}) can be read, from left to right, along the ‘W’ shape of the lower edge as G(ax^{∗}) ≅ Xd^{∗}.XaA.Xx^{∗}A.dA. But each of the squares in the diagram is a (product-absolute) pullback so that with Proposition 4.8 at hand we can continue:

Xd^{∗}.XaA.Xx^{∗}A.dA ≅ Xa.X(S, a)^{∗}.(x, S)A.x^{∗}A ≅ Xa.(x, S).(S, a)^{∗}.x^{∗}A ≅ (x, a)(x, a)^{∗}

as required.

5. Characterization of Bicategories of Spans

5.1. If B is a cartesian bicategory with MapB essentially locally discrete then each slice MapB/(X ⊗ A) is also essentially locally discrete and we can write Span MapB(X, A) for the categories obtained by taking the quotients of the equivalence relations comprising the hom categories of the MapB/(X ⊗ A). Then we can construct functors C_{X,A}:Span MapB(X, A) [figure] B(X, A), where for an arrow in Span MapB(X, A) as shown,

[figure: A; N; b; M; A; a; M; N; h; X; N; y; M; X; x; M; N; h]

we define C(y, N, b) = by^{∗} and C(h) : ax^{∗} = (bh)(yh)^{∗} ≅ bhh^{∗}y^{∗} [figure: bϵ_{h}y^{∗}] by^{∗}. If MapB is known to have pullbacks then the Span MapB(X, A) become the hom-categories for a bicategory Span MapB and we can consider whether the C_{X,A} provide the effects on homs for an identity-on-objects pseudofunctor C:Span MapB [figure] B. Consider

[figure: Y; A; N; Y; y; N; A; b; N; A; N; M; M; A; A; X; M; A; a; M; X; x; N; M; P; N; p; P; M; r]  (6)

where the square is a pullback. In somewhat abbreviated notation, what is needed further are coherent, invertible 2-cells C̃:CN.CM [figure] C(NM) = CP, for each composable pair of spans M, N, and coherent, invertible 2-cells C^{∘}:1_{A} [figure] C(1_{A}), for each object A. Since the identity span on A is (1_{A}, A, 1_{A}), and C(1_{A}) = 1_{A}.1_{A}^{∗} ≅ 1_{A}.1_{A} ≅ 1_{A} we take the inverse of this composite for C^{∘}. To give the C̃ though is to give 2-cells yb^{∗}ax^{∗} [figure] ypr^{∗}x^{∗} and since spans of the form (1_{N}, N, b) and (a, M, 1_{M}) arise as special cases, it is easy to verify that to give the C̃ it is necessary and sufficient to give coherent, invertible 2-cells b^{∗}a [figure] pr^{∗} for each pullback square in MapB. The inverse of such a 2-cell pr^{∗} [figure] b^{∗}a is the mate of a 2-cell bp [figure] aq. But by discreteness a 2-cell bp [figure] aq must be essentially an identity. Thus, definability of C̃ is equivalent to the invertibility in B of the mate pr^{∗} [figure] b^{∗}a of the identity bp [figure] ar, for each pullback square as displayed in (6). In short, if MapB has pullbacks and these satisfy the Beck condition as in Proposition 4.8 then we have a canonical pseudofunctor C:Span MapB [figure] B.

5.2. Theorem. For a bicategory B the following are equivalent:

i) There is a biequivalence B ≃ Span E, for E a category with finite limits;

ii) The bicategory B is cartesian, each comonad has an Eilenberg-Moore object, and every map is comonadic.

iii) The bicategory MapB is an essentially locally discrete bicategory with finite limits, satisfying in B the Beck condition for pullbacks of maps, and the canonical

C:Span MapB [figure] B

is a biequivalence of bicategories.

Proof. That i) implies ii) follows from our discussion in the Introduction. That iii) implies i) is trivial so we show that ii) implies iii).

We have already observed in Theorem 3.14 that, for B cartesian with every object discrete, MapB is essentially locally discrete and we have seen by Propositions 2.3 and 3.5 that, in a cartesian bicategory in which Maps are Comonadic, every object is discrete. In Theorem 4.8 we have seen that, for B satisfying the conditions of ii), MapB has pullbacks, and hence all finite limits and, in B the Beck condition holds for pullbacks. Therefore we have the canonical pseudofunctor C:Span MapB [figure] B developed in 5.1. To complete the proof it suffices to show that the C_{X,A}: Span MapB(X, A) [figure] B(X, A) are equivalences of categories.

Define functors F_{X,A}:B(X, A) [figure] Span MapB(X, A) by F (R) = F_{X,A}(R) = (u, τR, v) where

[figure: τR; A; v; X; τR; u; X; A; R; ω]

is the R-component of the counit for ι ⊣ τ:G [figure] MapB. For a 2-cell α:R [figure] R′ we define F (α) to be the essentially unique map satisfying

[figure: τR; τR′; F (α); τR′; A; v′; X; τR′; u′; X; A; R′; ω′; τR; X; u; τR; A; v; =; τR; A; v; X; τR; u; X; A; X; A; R; X; A; R′; ω; α]

(We remark that essential uniqueness here means that F (α) is determined to within unique invertible 2-cell.) Since ω:v [figure] Ru has mate vu^{∗} [figure] R invertible, because (v, τR, u) is a tabulation of R, it follows that we have a natural isomorphism CF R [figure] R. On the other hand, starting with a span (x, S, a) from X to A we have as a consequence of Theorem 4.9 that (x, S, a) is part of a tabulation of ax^{∗}:X [figure] A. It follows that we have a natural isomorphism (x, S, a) [figure] F C(x, S, a), which completes the demonstration that C_{X,A} and F_{X,A} are inverse equivalences.

6. Direct sums in bicategories of spans

In the previous section we gave a characterization of those (cartesian) bicategories of the form Span E for a category E with finite limits. In this final section we give a refinement, showing that Span E has direct sums if and only if the original category E is lextensive [CLW].

Direct sums are of course understood in the bicategorical sense. A zero object in a bicategory is an object which is both initial and terminal. In a bicategory with finite products and finite coproducts in which the initial object is also terminal there is a canonical induced arrow X +Y [figure] X ×Y, and we say that the bicategory has direct sums when this map is an equivalence.

6.1. Remark. Just as in the case of ordinary categories, the existence of direct sums gives rise to a calculus of matrices. A morphism X_{1} + … + X_{m} [figure] Y_{1} + … + Y_{n} can be represented by an m × n matrix of morphisms between the summands, and composition can be represented by matrix multiplication.

6.2. Theorem. Let E be a category with finite limits, and B = Span E. Then the following are equivalent:

i) B has direct sums;

ii) B has finite coproducts; iii) B has finite products;

iv) E is lextensive.

Proof. [i) ⟹ ii)] is trivial.

[ii) ⟺ iii)] follows from the fact that B^{op} is biequivalent to B.

[ii) ⟹ iv)] Suppose that B has finite coproducts, and write 0 for the initial object and + for the coproducts.

For every object X there is a unique span 0 [figure] D [figure] X. By uniqueness, any map into D must be invertible, and any two such with the same domain must be equal. Thus when we compose the span with its opposite, as in 0 [figure] D [figure] X [figure] D [figure] 0, the resulting span is just 0 [figure] D [figure] 0. Now by the universal property of 0 once again, this must just be 0 [figure] 0 [figure] 0, and so D ≅ 0, and our unique span 0 [figure] X is a map.

Clearly coproducts of maps are coproducts, and so the coproduct injections X + 0 [figure] X + Y and 0 + Y [figure] X + Y are also maps. Thus the coproducts in B will restrict to E provided that the codiagonal u:X + X [figure] E [figure] X:v is a map for all objects X. Now the fact that the codiagonal composed with the first injection i : X [figure] X + X is the identity tells us that we have a diagram as on the left below

\xymatrix{ &  & X \ar[dr]_{i′} \ar[ddrr]^{1} &  &  &  &  & X \ar[dr]_{i′} \ar[ddrr]^{i} \\  & X \ar[dr]^{i} &  & E \ar[dl]_{u} \ar[dr]_{v} &  &  & X \ar[dr]^{i} &  & E \ar[dl]_{u} \ar[dr]_{u} \\ X &  & X + X &  & X & X &  & X + X &  & X + X}

in which the square is a pullback; but then the diagram on the right shows that the composite of u:X+X [figure] E [figure] X+X:u with the injection i : X [figure] X+X is just i. Similarly its composite with the other injection j : X [figure] X+X is j, and so u:X+X [figure] E [figure] X+X:u is the identity. This proves that the codiagonal is indeed a map, and so that E has finite coproducts; we have already assumed that it has finite limits. To see that E is lextensive observe that we have equivalences

E/(X + Y) ≃ B(X + Y, 1) ≃ B(X, 1) × B(Y, 1) ≃ E/X × E/Y.

[iv) ⟹ i)] Suppose that E is lextensive. Then in particular, it is distributive, so that (X + Y) × Z ≅ X × Z + X × Y, and we have

B(X + Y, Z) ≃ E/((X + Y) × Z) ≃ E/(X × Z + Y × Z)
≃ E/(X × Z) × E/(Y × Z) ≃ B(X, Z) × B(Y, Z)

which shows that X + Y is the coproduct in B; but a similar argument shows that it is also the product.

6.3. Remark. The implication iv) ⇒ i) was proved in [P&S, Section 3].

6.4. Remark. The equivalence ii) ⇔ iv) can be seen as a special case of a more general result [H&S] characterizing colimits in E which are also (bicategorical) colimits in Span E.

6.5. Remark. There is a corresponding result involving partial maps in lextensive categories, although the situation there is more complicated as one does not have direct sums but only a weakened relationship between products and coproducts, and a similarly weakened calculus of matrices. See [C&L, Section 2].

There is also a nullary version of the theorem. We simply recall that an initial object in a category E is said to be strict, if any morphism into it is invertible, and then leave the proof to the reader. Once again the equivalence ii) ⇔ iv) is a special case of [H&S].

6.6. Theorem. Let E be a category with finite limits, and B = Span E. Then the following are equivalent:

i) B has a zero object;

ii) B has an initial object;

iii) B has a terminal object;

iv) E has a strict initial object.

References

[CKWW] A. Carboni, G.M. Kelly, R.F.C. Walters, and R.J. Wood. Cartesian bicategories II, Theory Appl. Categ. 19 (2008), 93–124.

[CLW] A. Carboni, Stephen Lack, and R.F.C. Walters. Introduction to extensive and distributive categories. J. Pure Appl. Algebra 84 (1993), 145–158.

[C&W] A. Carboni and R.F.C. Walters. Cartesian bicategories. I. J. Pure Appl. Algebra 49 (1987), 11–32.

[C&L] J.R.B. Cockett and Stephen Lack. Restriction categories III: colimits, partial limits, and extensivity, Math. Struct. in Comp. Science 17 (2007), 775–817.

[LSW] I. Franco Lopez, R. Street, and R.J. Wood, Duals Invert, Applied Categorical Structures, to appear.

[H&S] T. Heindel and P. Sobociński, Van Kampen colimits as bicolimits in Span, Lecture Notes in Computer Science, (CALCO 2009), 5728 (2009), 335–349.

[P&S] Elango Panchadcharam and Ross Street, Mackey functors on compact closed categories, J. Homotopy and Related Structures 2 (2007), 261–293.

[ST] R. Street. The formal theory of monads, J. Pure Appl. Algebra 2 (1972), 149–168.

[W&W] R.F.C. Walters and R.J. Wood. Frobenius objects in cartesian bicategories, Theory Appl. Categ. 20 (2008), 25–47.

School of Computing and Mathematics University of Western Sydney Locked Bag 1797 South Penrith DC NSW 1797 Australia and

Dipartimento di Scienze delle Cultura Politiche e dell’Informazione Università dell Insubria, Italy and Department of Mathematics and Statistics Dalhousie University Halifax, NS, B3H 3J5 Canada

Email: s.lack@uws.edu.au
      robert.walters@uninsubria.it
      rjwood@dal.ca

This article may be accessed at http://www.tac.mta.ca/tac/ or by anonymous ftp at

ftp://ftp.tac.mta.ca/pub/tac/html/volumes/24/1/24-01.{dvi,ps,pdf}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX2e strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis, cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown (at) btinternet.com Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it Valeria de Paiva, Cuill Inc.: valeria@cuill.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, University of Western Sydney: s.lack@uws.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Tom Leinster, University of Glasgow, T.Leinster@maths.gla.ac.uk Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Brooke Shipley, University of Illinois at Chicago: bshipley@math.uic.edu James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca